OpenStudy (anonymous):

. Which pair has equally likely outcomes? List the letters of the two choices below which have equal probabilities of success, separated by a comma. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards). A. drawing a black seven out of a standard 52 card deck given it’s not a face card or an Ace. B. rolling a sum of 4 on two fair six sided dice C. drawing a three out of a standard 52 card deck given it’s not a face card or an Ace. D. rolling a sum of 9 on two fair six sided dice E. rolling a sum of 7 on two fair six sided dice

OpenStudy (anonymous):

do you want the probabilities of each?

OpenStudy (anonymous):

working backwards, E has probability \[\frac{1}{6}\]

OpenStudy (anonymous):

well it just asks which pair has equally like out comes

OpenStudy (anonymous):

D has probability \[\frac{4}{36}=\frac{1}{9}\]

OpenStudy (anonymous):

there are 12 face cards and 4 aces so if you know you did not get one of those there are 52-16 = 36 cards left to choose from and 2 black sevens so the probability of A is \[\frac{2}{36}=\frac{1}{18}\]

OpenStudy (anonymous):

that was A sorry i am going out of order

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

rolling as sum of 4 on two dice has probability \[\frac{3}{36}=\frac{1}{12}\]

OpenStudy (anonymous):

i have yet to see the same number twice. maybe the last one?

OpenStudy (anonymous):

i dont think i have to give the probablily i think i just have to list the 2 that are..

OpenStudy (anonymous):

well in order to find them we need to compute. i haven't seen the same number twice.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

last one... there are 4 threes. not face card or ace means you have 36 cards left in the deck. so the probability you get a 3 given it is not a face card or an ace is \[\frac{4}{36}=\frac{1}{9}\] whew

OpenStudy (anonymous):

same as probability of rolling a 9 on two fair dice. what a pain!

OpenStudy (anonymous):

lol the last problems you helped me on.. helped get my grade up to a 60% intstead of 44

OpenStudy (anonymous):

still there?